# Weakly-supervised diagnosis identification from Italian discharge letters

### v1

@@ 2.0.2 Semi-automatic labelling pipeline (SAL)
2.0.3 Classification of discharge letters using the weakly-labelled dataset (WLC)
@@
The classification pipeline that exploits the weakly labelled data produced by Step 1 corresponds to Step 2 in Figure 1 (WLC). A pre-processing step (Step 2.1) is necessary to remove headers and footers that contain information that is non-relevant to our analysis. Supplementary Material B reports details about this. The classification model we adopted is based on the fine-tuned version of the Umberto model that we presented in Section 2.0.1 (BPT). In this case, the transformer-based model is enhanced with a linear classification layer that outputs the probability of having the disease. It is trained with the discharge letters’ text and the weak labels assigned to the letters based on the clusters to which their diagnoses have been assigned (Step 2.2).
@@
Given the limitation of BERT-based models in dealing with texts longer than 512 tokens, when this is the case, we truncated the letter text to the first 512 tokens after pre-processing, possibly including the diagnosis sentences, even if at the end of the letter (in the setting in which we kept it). This is not a relevant limitation, considering that 90% of discharge letters in our dataset are within 512 tokens after pre-processing. In the case of datasets with longer texts, it is possible to chunk them in portions no longer than 512 tokens, process them in parallel and combine their outputs, taking the maximum probability.
2.1 Training and evaluation
The BERT-based classifier is trained with the AdamW optimizer, using a linear learning rate scheduler starting from $10^{-3}$, with a weight decay of 0.01, for six epochs. Initially, only the last classification layer is trained, and after two epochs, the last BERT layer is unfrozen. After four epochs, the last but one BERT layer is unfrozen. These hyperparameters have been selected with grid search.
We cannot perform a precise separate evaluation of all the pipeline steps since we do not have ground truth labels related to diagnosis strings and all their clusters. Because of this, we assess the performance of the whole pipeline on the classification task, implicitly evaluating all the previous steps. For the clustering, we can perform a specific evaluation limited to the clusters related to our bronchiolitis gold labels.
[…]
The classification results are evaluated in terms of precision, recall and F1-score for the positive class and area under the receiver operating characteristic curve (AUC). These metrics are computed for both our weak labels and the gold standard bronchiolitis labels. Evaluating against weak labels allows for an assessment of the classification model’s performance alone, while evaluation against gold standard labels provides a comprehensive view of the pipeline’s effectiveness. We exclude AUC for rule-based classifiers, as they do not produce probability estimates, and we do not assess the fully supervised classifier using weak labels. We compute the metrics with 10-fold stratified cross-validation.
@@ 2.1 Training and evaluation
    Removing the diagnosis strings exploited to assign weak labels
Since our dataset contains data from different hospitals and different LHUs, we verify the presence of differences among them. Since the number of cases for each hospital and LHU is limited, we consider those with a minimum number of bronchiolitis cases (15) and measure the results using a Leave One Group Out (LOGO) approach. For each hospital/LHU to be evaluated, we train the model on the data from the other hospitals/LHUs, and we measure the performance on the data from the hospital/LHU of interest.
We also compare results between the groups of pediatric vs non-pediatric ER/Department admissions. In this case, we adopt the same 10-fold stratified cross-validation used to compare different classification models, dividing the test set between pediatric and non-pediatric ERs/Departments at each iteration and computing the metrics on these two groups separately.
@@ 3 Results
In this section, we present the results of our pipeline for the bronchiolitis case study, for the different stages that compose it. In particular, Section 3.1 describes the dataset, Section 3.2 presents results related to the extraction of diagnosis strings from discharge letters, Section 3.3, covers the clustering pipeline and Section 3.4 presents the final results on the identification of bronchiolitis.
@@ 3.1 Dataset
The dataset comprises 33’176 discharge letters related to children who were admitted to Emergency Rooms and/or hospitalized in the Veneto Region between 2017 and 2020. The dataset covers 44 different hospitals in the region, which can be grouped into 9 Local Health Units (LHUs) and 2 university hospitals. Each discharge letter in the dataset includes the complete text of the letter, along with the date of the letter and anonymized identifiers of the patient. A manual review of discharge letters was conducted to extract relevant diagnoses of bronchiolitis, resulting in 335 positive cases within the dataset. This review was performed by a data manager with a clinical background, with the support of clinicians for uncertain cases. These labels are the gold standard to validate our pipeline, and we do not make direct use of them in the pipeline training.
We identified the records related to pediatric ERs/Departments by looking at the presence of specific keywords (e.g., pediatria, pediatrico, pediatrica) in the header of the letters. This marked 10’340 records (31.17%) as related to pediatric ERs/Departments. Supplementary Material A1 reports additional plots with data distributions among hospitals and LHUs. In particular, ten hospitals cover more than 85% of the records.
The median length of the letters is 4515 characters (1st-3rd quartiles 3240-5806). An example of a fictitious letter is reported in Supplementary Material A2.
@@ 3.2 Diagnosis extraction
Our diagnosis extraction module (Step 1.1) successfully extracted a diagnosis sentence from 89% of the discharge letters in the dataset. We extracted 15’038 unique diagnosis sentences, of which 75% occur only once, confirming the need for clustering them in wider groups that can be more easily related to known diseases and symptoms. A few of the extracted strings include the ICD9-CM codes and their standard textual description. Specifically, 737 diagnosis strings (4.9%) contain an ICD9-CM code, covering 3’436 discharge letters (10.4%). The most common strings are the simplest, while more articulate ones have single occurrences. The median length of the extracted diagnoses strings is 49 characters, with 1st and 3rd quartiles of 17 and 156 characters. Among the 335 bronchiolitis-labelled letters, we extracted 138 unique diagnosis strings from 306 letters.
@@ 3.3 Diagnosis clustering
In the first clustering level, the 15’038 extracted diagnosis strings were grouped into 590 clusters. Cluster sizes varied significantly, from a large cluster of 270 strings to several smaller clusters containing 5 strings. Table 1 presents the 20 largest clusters, including their sizes and descriptive keywords. The subsequent second-level clustering reduced the number of clusters to 135.
@@
For our case study on bronchiolitis identification, Table 2 details the main clusters related to respiratory infections, including the number of bronchiolitis cases they cover. Four clusters (#1, #2, #3, #5) contain the keyword “bronchiolitis”. Three include a large majority of bronchiolitis cases, whereas the fourth cluster (#5) has only half of its diagnoses labelled as bronchiolitis. These four clusters together cover 65.38% of all bronchiolitis cases. Additionally, 13 clusters include the keyword “bronchospasm”. According to physicians who annotated the data, bronchospasm is associated with bronchiolitis only if accompanied by fever. Only one of these 13 clusters (#4) has both “bronchospasm” and “fever” keywords, and indeed, it contains a high percentage (70%) of bronchiolitis, accounting for an additional 15% of total bronchiolitis cases. Thus, these five clusters are those that, without knowing the gold labels, would be selected to define the weak labels for the training set, according to the two definitions of {bronchiolitis} and {bronchospasm, fever}. Among the respiratory infection clusters, three other clusters contain bronchiolitis cases (#6, #7, #8), but they are in a limited number.
@@
For the second-level clustering, Table 3 presents clusters related to respiratory infections. Some clusters from Table 2 have been merged: two related to mild bronchiolitis, eight related to bronchospasm, and five about generic respiratory infections. The selection of weak labels for training remains unchanged since, looking at their keywords, we select second-level clusters #1, #2, #3, #4 that cover the same records as the clusters we would have chosen at the first level. The main advantage of second-level clustering is the substantial reduction in the number of clusters (approximately 1/5 of the original number of clusters in our dataset), simplifying weak labelling, especially if manual verification is involved.
@@
This cluster selection for weak labels covers 80% of the original bronchiolitis cases. To understand the reason for the false negatives, we reviewed twenty letters labelled as bronchiolitis in the gold standard but not covered by our weak labels. We found 10 cases where the diagnosis pertains to other symptoms or disease (e.g., otitis, dyspnea, fever), but the bronchiolitis is mentioned in the letter, at least as a suspect condition; 4 cases where bronchiolitis is mentioned in the diagnosis alongside other diseases, leading to assignment to a cluster related to the other disease; 3 cases of respiratory issues with no explicit mention of bronchiolitis, respiratory syncytial virus, or bronchospasm and fever; 2 cases with no diagnosis string present; 1 case where a family member has bronchiolitis, but the patient’s diagnosis is of a different disease. These results suggest that the absence of some gold labels in our weak labels primarily arises from the lack of explicit bronchiolitis indications, with only a few cases that could have been detected but were missed. This highlights the importance of using the full discharge letters for a model aiming to identify diseases accurately. Furthermore, there might be a few errors in the gold labels themselves.

### v2

@@ 2.2.3 Mapping clusters to diseases
2.3 Classification of discharge letters using the weakly-labelled dataset (WLC)
@@
The classification pipeline that exploits the weakly-labelled data produced by Step 1 corresponds to Step 2 (WLC) in Figure 1.a. A pre-processing step (Step 2.1) removes headers and footers that contain information that is non-relevant to our analysis (see B for details and results without pre-processing). The classification model we adopted is based on the fine-tuned version of the UmBERTo model that we presented in Section 2.1. In this case, the transformer-based model is enhanced with a linear classification layer that outputs the probability of having the disease. It is trained with the discharge letters’ text and the weak labels assigned to the letters based on the clusters to which their diagnoses have been assigned (Step 2.2).
@@
Given the limitation of RoBERTa-based models in dealing with texts longer than 512 tokens, when this is the case, we truncated the letter text to the first 512 tokens after pre-processing, possibly including the diagnosis sentences, even if at the end of the letter (in the setting in which we kept it). This is not a relevant limitation, considering that 90% of discharge letters in our dataset are within 512 tokens after pre-processing. In the case of datasets with longer texts, it is possible to chunk them in portions no longer than 512 tokens, process them in parallel and combine their outputs, taking the maximum probability.
2.4 Experimental outline
Our main evaluation is at the final classification stage of the pipeline, where we can compare the results against manually annotated labels. In addition to this, we evaluated the intermediate steps of the pipeline, in particular the diagnosis extraction (Step 1.1) and clustering (Steps 1.2-1.3), on a randomly selected subset of 1,000 diagnosis strings that have been manually annotated. Each of these strings was marked as correct, partially correct , or wrong with respect to the fact of being a complete diagnosis string and to the assignments to the first level and the second level clusters. Additional results on the clustering steps, including robustness analyses with respect to hyperparameter choices and clustering algorithms, are reported in D.
In evaluating classification results, we compare our pipeline with some fully unsupervised baselines: two rule-based approaches that do not require supervision and with a zero-shot LLM for the Italian language (LLaMAntino-3-ANITA-8B-Inst-DPO-ITA[polignano2024advanced]. The two rule-based approaches look for the presence of the terms corresponding to the definitions of bronchiolitis used in Step 1.3, with the first one (Rule-based full-text) looking into the entire text of the letter and the second one (Rule-based diagnosis) looking only in the diagnosis strings extracted by Step 1.1. The zero-shot LLM is prompted to indicate if the text of the text discharge letter indicates a diagnosis of bronchiolitis, answering with a “Si” (“Yes”) or a “No”.
Additionally, we assess in detail the performance of the classification model of our pipeline (Step 2.2) by replacing it with other classifiers:
  1. 1.
    The original UmBERTo model [tamburini2020bertology], without TPT (WS-UmBERTo)
  2. 2.
    XLM-RoBERTa-base [conneau2020unsupervised], a multilingual version of RoBERTa [liu2019roberta], with (WS-XLM-RoBERTa-TPT) and without (WS-XLM-RoBERTa) TPT
  3. 3.
    RemBERT [chung2020rethinking], an improved version of the multilingual mBERT, with (WS-RemBERT-TPT) and without (WS-RemBERT-BPT) without TPT
  4. 4.
    An LSTM-based recurrent neural network exploiting Italian Word2Vec embeddings [mikolov2013efficient] (WS-LSTM)
The three transformer-based models with pretraining are also evaluated in a fully supervised setting, training them on the manually annotated labels (FS-UmBERTo-TPT, FS-XLM-RoBERTa-TPT, FS-RemBERT-TPT), so as to have a measure of the performance loss due to weak supervision.
@@ 2.4 Experimental outline
    Removing the diagnosis strings exploited to assign weak labels (when present)
We also measure the performance of the models on the weak labels, to assess the difference with the ones observed on gold labels, and we verify the robustness of the results on gold labels to the weak label selection, measuring the effect of removing single clusters from positive weak labels. To further analyze the noise introduced by weak labels, we also experiment with replacing a varying percentage of gold labels with weak labels.
The classification results are evaluated in terms of precision (P), recall (R) and F1-score (F1) for the bronchiolitis class and area under the receiver operating characteristic curve (AUC). For classifiers producing a continuous score, the classification threshold is set via Youden’s index [youden1950index]. We exclude AUC for rule-based classifiers, as they do not produce probability estimates, and we do not assess the fully supervised models using weak labels. The positive class probability for the zero-shot LLM is estimated by applying a softmax function over the token-level log-likelihoods of the tokens “Si” (“Yes”) and “No”, following previous literature [brown2020language]. All metrics are computed with 10-fold stratified cross-validation.
Since our data come from different hospitals and different LHUs, we also perform additional experiments, reported in E, to verify the presence of differences among them.
2.5 Training details
The additional pre-training of the transformer-based models over the E3C corpus was performed with the masked-word prediction task, with a learning rate of $2\cdot 10^{-5}$ and weight decay of $0.01$, using 20% of the data as a validation set for early stopping.
The LSTM baseline has the frozen Word2Vec embedding layer followed by a 32-unit LSTM layer, a dropout layer, a 16-unit fully connected layer, a dropout layer and the final output layer. It is trained with Adam optimizer, with a learning rate scheduler starting from $10^{-2}$, a decay rate of 0.7, 100 decay steps, using early stopping.
The transformer-based classifiers are trained with the AdamW optimizer, using a linear learning rate scheduler starting from $10^{-3}$, with a weight decay of 0.01, for a number of epochs regulated by early stopping (leading to 6 epochs for the UmBERTo models, 4 for the XLM-RoBERTa models and 3 for the RemBERT models). Initially, only the last classification layer is trained, and after two epochs, the last transformer layer is unfrozen. After four epochs, if reached, the last but one transformer layer is unfrozen. These hyperparameters have been selected with grid search.
Due to the class imbalance, in the training folds the negative class is undersampled, limiting its size to twice that of the positive class.
@@ 3.1 Dataset
The dataset for the case study comprises 33,176 discharge letters related to children who were admitted to Emergency Rooms and/or hospitalized in the Veneto Region between 2017 and 2020, with a family pediatrician adhering to the Pedianet network [cantarutti2021pedianet]. Inclusion in the Pedianet database is voluntary and parents/legal guardians must provide informed consent for their children’s anonymized data to be used for research purposes. The study was performed in accordance with the Declaration of Helsinki. Ethical approval of the study and access to the database were granted by the Internal Scientific Committee of So.Se.Te. Srl, the legal owner of Pedianet.
The dataset covers 44 different hospitals in the region, which can be grouped into 9 Local Health Units (LHUs) and 2 university hospitals. Each discharge letter includes the complete text of the letter, along with the date of the letter and an anonymized identifier of the patient. A manual review of discharge letters was conducted to extract relevant diagnoses of bronchiolitis, resulting in 335 positive cases within the dataset. This review was performed by a data manager with a clinical background, with the support of clinicians for uncertain cases. These labels are the gold standard to validate our pipeline, and we do not make direct use of them in the pipeline training (except for the fully-supervised experiments).
A.1 reports additional plots with data distributions among hospitals and LHUs. In particular, 10 hospitals cover more than 85% of the records.
The median length of the letters is 4,515 characters (1st-3rd quartiles 3,240-5,806). An example of a fictitious letter is reported in A.2.
@@ 3.2 Diagnosis extraction
Our diagnosis extraction module (Step 1.1) successfully extracted a diagnosis sentence from 89% of the discharge letters in the dataset. We extracted 15,038 unique diagnosis sentences, of which 75% occurred only once, confirming the need for clustering them in wider groups that can be more easily related to known diseases and symptoms. ICD-9-CM codes are present only in 737 diagnosis strings (4.9%), covering 3,436 discharge letters (10.4%). The most common strings are the simplest, while more articulate ones have single occurrences. The median length of the extracted diagnoses strings is 49 characters, with first and third quartiles of 17 and 156 characters. Among the 335 bronchiolitis-labelled letters, we extracted 138 unique diagnosis strings from 306 letters. The manual validation of a sample of 1,000 extracted diagnosis strings revealed that 985 (98.5%) were correct diagnosis strings, only 4 (0.4%) were not diagnosis strings, and 11 (1.1%) were incomplete diagnosis strings.
@@ 3.3 Diagnosis clustering
In the first clustering level, the 15,038 extracted diagnosis strings were grouped into 590 clusters. Cluster sizes varied significantly, from a large cluster of 270 strings to several smaller clusters containing 5 strings. The subsequent second-level clustering reduced the number of clusters to 135. Summary of main clusters is reported in D.
Table 1: Manual evaluation of the clustering assignment for 1,000 diagnosis strings. comparing results after first level and second level clustering.
Correct Partially Correct Wrong
1st level clustering 69.60 % 22.20 % 8.20 %
2nd level clustering 71.90 % 20.30 % 7.80 %
$\Delta$ 2nd-1st 2.30 % -1.90 % -0.40 %
Table 1 summarizes the manual evaluation of the clustering assignment, at both levels, for the annotated sample of 1,000 strings. Most of the partially correct cluster assignments are due to details of the diseases, e.g., fracture of right hand assigned to a cluster for fracture of left hand, severe influenza assigned to a cluster for mild influenza. A few cases are related to multiple symptoms/diseases not completely matching the cluster, such as vomiting and fever in the cluster for fever alone.
@@
For our case study on bronchiolitis identification, Table 2 details the main clusters related to respiratory infections, including the number of bronchiolitis cases they cover. Given space constraints, we report here only clusters mentioning bronchiolitis, bronchospasm, pneumonia, or respiratory infections, as these are the clusters most closely related to bronchiolitis or its associated symptoms. Four clusters (#1, #2, #3, #5) contain the keyword “bronchiolitis”. Three of them include a large majority of bronchiolitis cases, whereas the fourth cluster (#5) has only half of its diagnoses labelled as bronchiolitis. These four clusters together cover 65.38% of all bronchiolitis cases. Additionally, 13 clusters include the keyword “bronchospasm”, a symptom frequently associated with bronchiolitis. According to physicians who annotated the data, bronchospasm is associated with bronchiolitis only if accompanied by fever. Only one of these 13 clusters (#4) has both “bronchospasm” and “fever” keywords, and indeed, it contains a high percentage (70%) of bronchiolitis, accounting for an additional 15% of total bronchiolitis cases. Thus, these five clusters are those that, without knowing the gold labels, would be selected to define the weak labels for the training set, according to the two keyword-based definitions of {positive={bronchiolitis},negative={}} and {positive={bronchospasm, fever},negative={}}. Among the respiratory infection clusters, three other clusters contain bronchiolitis cases (#6, #7, #8), but they are in a limited number.
@@
For the second-level clustering, Table 3 presents clusters related to respiratory infections. The second level clustering appears to have effectively merged some of the clusters that were previously separated in Table 2: two related to mild bronchiolitis, eight related to bronchospasm, and five about generic respiratory infections. The selection of weak labels for training remains unchanged since, looking at their keywords, we select second-level clusters #1, #2, #3, #4 that cover the same records as the clusters we would have chosen at the first level. Second-level clustering determines a substantial reduction in the number of clusters (approximately 1/5 of the original number of clusters in our dataset), simplifying manual inspection and determining a positive effect on diagnosis-cluster mappings as observed in Table 1.
@@
This cluster selection for weak labels covers 80% of the original bronchiolitis cases. To understand the reason for the false negatives, we reviewed 20 letters labelled as bronchiolitis in the gold standard but not covered by our weak labels. We found 10 cases where the diagnosis pertains to other symptoms or disease (e.g., otitis, dyspnea, fever), but the bronchiolitis is mentioned in the letter, at least as a suspect condition; 4 cases where bronchiolitis is mentioned in the diagnosis alongside other diseases, leading to assignment to a cluster related to the other disease; 3 cases of respiratory issues with no explicit mention of bronchiolitis, respiratory syncytial virus, or bronchospasm and fever; 2 cases with no diagnosis string present; 1 case where a family member has bronchiolitis, but the patient’s diagnosis is of a different disease. These results suggest that the absence of some gold labels in our weak labels primarily arises from the lack of explicit bronchiolitis indications, with only a few cases that could have been detected but were missed. They also highlight the importance of using the full discharge letters for a model aiming to identify diseases accurately.